OpenStudy (anonymous):

Write a proportion for the following. When surveyed, 5 out of every 8 students liked the food in the cafeteria. If 60 liked the food, how many were surveyed? A. 5/8 = x/60 B. 5/8 = 60/x C. 5/60 = x/8 I think its C but i a not sure

OpenStudy (anonymous):

B.

OpenStudy (ranga):

Proportion here is: Student who liked the food / Total number of students = 5 / 8 = 60 / x

OpenStudy (anonymous):

dorian explain please

OpenStudy (anonymous):

Okay. 5 our of every 8 liked it, right? so you could right that as 5/8. Now, you know that 60 students liked it, but you don't know how many didn't like it. So, you know that 60 equals the 5 part, because they are both the people who liked the food. However, you do not know what number equals the 8 part, who didn't like the food, so therefore 5/8 = 60/x.

OpenStudy (ranga):

^^^ Not who did not like the food but x and 8 are the total surveyed.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Oops - my bad on what y was.

OpenStudy (anonymous):

*write*

OpenStudy (anonymous):

I don't get it